OpenStudy (anonymous):

John and George together raked 7/8 of the yard.John raked 3/4 of that amount.What part of the yard did John rake?

OpenStudy (anonymous):

so both John and George together do 7/8 of the yard. John does 3/4 of that 7/8 of the yard. How much does John do?

OpenStudy (anonymous):

3/4 * 7/8 ~

OpenStudy (anonymous):

\[3/4\times7/8\]

OpenStudy (anonymous):

answer is ?

OpenStudy (anonymous):

\[\frac{ a }{ b }*\frac{ c }{ d }=\frac{ ac }{ bd }\]

OpenStudy (anonymous):

\[\frac{3}{4} \times \frac{7}{8} = \frac{21}{32}\]